# Semidiscrete optimal transport with unknown costs

### v3

@@ 5.1 Properties of the objective
Let Assumptions 3.1-3.2 hold, and take $\delta\in\mathds{R}_{K}$ satisfying $\delta_{1}=0$. Fix $b_{1},...,b_{K}$ and let $\Delta_{k}=b_{k}-\beta_{k}$. Define $\Delta^{\max}=\max_{k}\|\Delta_{k}\|_{2}$. The following inequality holds:
@@
$\geq$ $\frac{\kappa_{2}}{8K}\left(\sum^{K}_{k=1}\min\left\{\delta^{2}_{k},\kappa^{2}_{1}\right\}1_{\left\{\left|\delta_{k}\right|\geq 8\kappa_{4}\Delta^{\max}\right\}}\right)1_{\left\{\kappa_{1}\geq 8\kappa_{4}\Delta^{\max}\right\}}.$
@@
$\delta^{\top}\left(p^{b,g^{*}+\delta}-p^{b,g^{*}}\right)\geq\max\left\{0,C\min\left\{\|\delta\|^{2}_{2},\kappa^{2}_{1}\right\}-C^{\prime}\left(\Delta^{\max}\right)^{2}\right\}$
@@
Consider the case $\Delta^{\max}\leq\frac{\kappa_{1}}{2\kappa_{4}}$. By Assumption 3.2(i), (25) yields
$P\left(\left|Z^{\beta}_{j,k}\right|<2\kappa_{4}\Delta^{\max}\right)\leq 4\kappa_{3}\kappa_{4}\Delta^{\max}.$
@@
$\left|p^{b,g^{*}}_{k}-p_{k}\right|\leq 4K\kappa_{3}\kappa_{4}\Delta^{\max}\cdot 1_{\left\{\Delta^{\max}\leq\frac{\kappa_{1}}{2\kappa_{4}}\right\}}+2\cdot 1_{\left\{\Delta^{\max}>\frac{\kappa_{1}}{2\kappa_{4}}\right\}}.$ (26)
@@ 5.2 Proof outline for main rate results
Let Assumptions 3.1-3.2 hold. For arbitrary $b,g$, let $\delta=g-g^{*}$ and $\Delta^{\max}=\max_{j}\|b_{j}-\beta_{j}\|_{2}$. Define $\pi^{b,g}\left(X\right)=\arg\min_{j}b^{\top}X_{j}-g_{j}$. Then, there exists $C_{0}>0$ such that
@@
Let Assumptions 3.1-3.2 and 3.3 hold. For arbitrary $b,g$, let $\delta=g-g^{*}$ and $\Delta^{\max}=\max_{j}\|b_{j}-\beta_{j}\|_{2}$. Define $\pi^{b,g}\left(X\right)=\arg\min_{j}b^{\top}X_{j}-g_{j}$, and let
@@
Let Assumptions 3.1-3.2 and 3.3 hold. Let $T$ be a deterministic subset of $\left\{1,...,n\right\}$, and define $V_{T}=\sum_{n\in T}X^{n}\left(X^{n}\right)^{\top}$. Then, for any $z>0$,
@@
Let Assumptions 3.1-3.2 and 3.3 hold. Let $\left\{\pi^{n}\right\}^{\infty}_{n=1}$ be a sequence of random variables such that each $\pi^{n}$ takes values in $\left\{1,...,K\right\}$ and is measurable with respect to ¯$\bar{\mathcal{F}}^{n}$, the sigma-algebra generated by $\pi^{1},...,\pi^{n-1}$, $W^{1}_{\pi^{1}},...,W^{n-1}_{\pi^{n-1}}$, and $X^{1},...,X^{n}$.
@@
Let Assumptions 3.1-3.2 and 3.3 hold, and let $\pi$ be the semi-myopic policy of Section 3.2. There exist constants $C_{3},C_{4},C_{5}>0$, which depend only on $d$, $K$ and $\kappa_{4},\kappa_{5},\kappa_{6}$, such that
@@
Let Assumptions 3.1-3.2 and 3.3 hold. Suppose that the semi-myopic policy of Section 3.2 is used to sample cost functions, while $g^{n}$ is updated as in Algorithm 1. Let $C_{1}$ be the constant obtained from Lemma 5.5. There exists a constant $C_{6}>0$ such that
@@
Let Assumptions 3.1-3.2 and 3.3 hold, and let $\pi$ be the semi-myopic policy of Section 3.2. There exists a constant $C_{7}>0$ such that
@@
Let Assumptions 3.1-3.2 and 3.3 hold. Suppose that the semi-myopic policy of Section 3.2 is used to sample cost functions, while $g^{n}$ is updated as in Algorithm 1. There exists some $\bar{\alpha}>0$ such that
@@
$\leq$ $\|\delta^{n}\|^{2}_{2}+\frac{2\alpha^{2}}{\left(n+1\right)^{2}}-\frac{2\alpha}{n+1}\left(\zeta^{n+1}-p^{\hat{\beta}^{n},g^{n}}\right)^{\top}\delta^{n}-\frac{2\alpha D_{1}}{n+1}\min\left\{\|\delta\|^{2}_{2},\kappa^{2}_{1}\right\}$
@@
$\mathbb{E}\left(\|\delta^{n+1}\|^{2}_{2}\mid\mathcal{F}^{n}\right)$$\leq$ $\|\delta^{n}\|^{2}_{2}+\frac{2\alpha^{2}}{\left(n+1\right)^{2}}-\frac{2\alpha D_{1}}{n+1}\min\left\{\|\delta\|^{2}_{2},\kappa^{2}_{1}\right\}$
@@
$\mathbb{E}\left(\|\delta^{n+1}\|^{2}_{2}\mid\mathcal{F}^{n}\right)$$\leq$ $\left(1-\frac{2\alpha D_{1}}{n+1}\min\left\{1,\frac{\kappa^{2}_{1}}{D^{2}_{0}}\right\}\right)\|\delta^{n}\|^{2}_{2}+\frac{2\alpha^{2}}{\left(n+1\right)^{2}}$
@@
$\mathbb{E}\left(\|\delta^{n+1}\|^{2}_{2}\mid\mathcal{F}^{n}\right)$$\leq$ $\left(1-\frac{2\alpha D_{1}}{n+1}\min\left\{1,\frac{\kappa^{2}_{1}}{D^{2}_{0}}\right\}\right)\|\delta^{n}\|^{2}_{2}+\frac{2\alpha^{2}}{\left(n+1\right)^{2}}$ (34)
@@
$\leq$ $\left(1-\frac{2\alpha D_{1}}{n+1}\min\left\{1,\frac{\kappa^{2}_{1}}{D^{2}_{0}}\right\}\right)\|\delta^{n}\|^{2}_{2}+\frac{2\alpha^{2}}{\left(n+1\right)^{2}}$
@@
$\mathbb{E}\left(\|\delta^{n+1}\|^{2}_{2}\right)$$\leq$ $\left(1-\frac{2\alpha D_{1}}{n+1}\min\left\{1,\frac{\kappa^{2}_{1}}{D^{2}_{0}}\right\}\right)\mathbb{E}\left(\|\delta^{n}\|^{2}_{2}\right)+\frac{2\alpha^{2}}{\left(n+1\right)^{2}}$
@@
$\leq$ $\left(1-\frac{2\alpha D_{1}}{n+1}\min\left\{1,\frac{\kappa^{2}_{1}}{D^{2}_{0}}\right\}\right)\mathbb{E}\left(\|\delta^{n}\|^{2}_{2}\right)+\frac{2\alpha^{2}}{\left(n+1\right)^{2}}$
@@
$\leq$ $\left(1-\frac{2\alpha D_{1}}{n+1}\min\left\{1,\frac{\kappa^{2}_{1}}{D^{2}_{0}}\right\}\right)\mathbb{E}\left(\|\delta^{n}\|^{2}_{2}\right)+\frac{2\alpha^{2}}{\left(n+1\right)^{2}}$
@@ 6 Numerical illustrations
We illustrate our framework on two numerical examples with linear costs. Section 6.1 considers synthetic test instances with Gaussian noise. Section 6.2 applies our framework to learn facility locations in a geographical partitioning problem (a common application area of semidiscrete optimal transport).
@@ 6.1 Synthetic linear costs
We consider a set of $1000$ synthetic test problems. In all of these problems, the data $X$ are generated uniformly on the unit sphere in $d=10$ dimensions, and the residual error follows a Gaussian distribution. Since $X$ can be represented as a normalized vector of independent standard Gaussian random variables, it is possible to verify Assumptions 3.1-3.3. Each problem uses $K=5$ different vectors $\beta_{k}$, also generated uniformly on the unit sphere, and a different $p$ generated uniformly on the $d$-dimensional simplex.
@@
For ease of implementation, we modified the semi-myopic policy to conduct forced exploration probabilistically: that is, in the $n$th iteration, we select $k$ uniformly at random with probability $\frac{\left(\frac{1}{a}\log n\right)^{9}}{n}$, where we used $a=4.5$. The rest of the time, we select a cost function according to the myopic policy $\hat{\pi}^{n}$. We used $\alpha=50$ in the update of $g^{n}$, and we set $\rho^{n}\equiv 0.001$ for simplicity, and in order to enable use of recursive least squares. At each $n$, we evaluate the policy using the correct selection indicator $1_{\left\{\hat{\pi}^{n}=\pi^{*}\left(X^{n}\right)\right\}}$. By averaging this quantity across $10$ runs, each containing $1000$ iterations and starting from $g^{0}\equiv 0$, $\hat{\beta}^{0}_{k}\equiv 0$, we obtain a trajectory of the probability $P\left(\hat{\pi}^{n}=\pi^{*}\left(X^{n}\right)\right)$ of correct selection. We then average this trajectory over $1000$ test problems.
@@
The optimal policy $\pi^{*}$ for each test problem is precomputed via brute force by running (12) with the true cost coefficients for a very large number of iterations. As a benchmark, we also implement a policy that knows the true costs (and, therefore, does not need to conduct any exploration or collect any samples), but not $g^{*}$, and runs (12) in an online manner. We also calculate the average probability of correct selection for this policy, which allows us to evaluate any loss incurred by having to learn the coefficients.
@@
Figure 1: Average performance of semi-myopic policy on synthetic problems.
@@
Figure 1 shows average performance for the semi-myopic policy under two levels of noise: $Var\left(\varepsilon\right)=0.02^{2}$ and $Var\left(\varepsilon\right)=0.2^{2}$. Since the benchmark policy knows the costs, it is unaffected by the noise variance. We see that, when the variance is low, there is virtually no performance loss caused by unknown costs. When the variance is high, the average PCS of the semi-myopic policy lags behind that of the known-cost policy by $0.022$ (and has narrowed this gap relative to the earlier iterations). Performance also does not appear to be very sensitive to the frequency of forced exploration: we also tested a much higher value $a=6$, for which the probability of forced exploration is only $0.0036$ by the end of the time horizon, but obtained very similar trajectories. This also indicates that, in an online implementation, we could allow forced exploration to happen quite infrequently while maintaining good performance.
Learning $g^{*}$ appears to be the greater challenge. After $1000$ iterations, the PCS is approximately $0.85$ even for the benchmark policy, which knows $\beta$ from the beginning. It is well-known in the stochastic approximation literature that the performance of SA can be sensitive to the stepsize rule. It may be possible to speed up convergence by using other choices of $\alpha_{n}$, many of which are surveyed in George & Powell (2006). We lightly tuned $\alpha$ in our stepsize rule $\alpha_{n}=\frac{\alpha}{n+1}$, but we do not delve deeply into this issue here because our focus in this paper is mainly theoretical.

### v4

@@ 5.1 Properties of the objective
Let Assumption 3.1 hold, and take $\delta\in\mathds{R}_{K}$ satisfying $\delta_{1}=0$. Fix $b_{1},...,b_{K}$ and let $\Delta_{k}=b_{k}-\beta_{k}$. Define $\Delta^{\max}=\max_{k}\|\Delta_{k}\|_{2}$. The following inequality holds:
@@
$\geq$ $\frac{\kappa^{\prime}_{2}\kappa^{\prime}_{4}}{8K}\left(\sum^{K}_{k=1}\min\left\{\delta^{2}_{k},\left(\kappa^{\prime}_{1}\right)^{2}\right\}1_{\left\{\min\left\{\left|\delta_{k}\right|,\kappa^{\prime}_{1}\right\}\geq 8\max\left\{1,\frac{\kappa^{\prime}_{3}}{\kappa^{\prime}_{2}\kappa^{\prime}_{4}}\right\}\kappa_{4}\Delta^{\max}\right\}}\right)1_{\left\{\Delta^{\max}\leq\frac{\kappa^{\prime}_{1}}{2\kappa_{4}}\right\}}.$
@@
$\delta^{\top}\left(p^{b,g^{*}+\delta}-p^{b,g^{*}}\right)\geq\max\left\{0,C\min\left\{\|\delta\|^{2}_{2},\left(\kappa^{\prime}_{1}\right)^{2}\right\}-C^{\prime}\left(\Delta^{\max}\right)^{2}\right\}$
@@
Consider the case $\Delta^{\max}\leq\frac{\kappa^{\prime}_{1}}{2\kappa_{4}}$. By Lemma 3.1(i), (34) yields
$P\left(\left|Z^{\beta}_{j,k}\right|<2\kappa_{4}\Delta^{\max}\right)\leq 4\kappa^{\prime}_{3}\kappa_{4}\Delta^{\max}.$
@@
$\left|p^{b,g^{*}}_{k}-p_{k}\right|\leq 4K\kappa^{\prime}_{3}\kappa_{4}\Delta^{\max}\cdot 1_{\left\{\Delta^{\max}\leq\frac{\kappa^{\prime}_{1}}{2\kappa_{4}}\right\}}+2\cdot 1_{\left\{\Delta^{\max}>\frac{\kappa^{\prime}_{1}}{2\kappa_{4}}\right\}}.$ (35)
@@ 5.2 Proof outline for main rate results
Let Assumption 3.1 hold. For arbitrary $b,g$, let $\delta=g-g^{*}$ and $\Delta^{\max}=\max_{j}\|b_{j}-\beta_{j}\|_{2}$. Define $\pi^{b,g}\left(X\right)=\arg\min_{j}b^{\top}X_{j}-g_{j}$. Then, there exists $C_{0}>0$ such that
@@
Let Assumption 3.1 and 3.2 hold. For arbitrary $b,g$, let $\delta=g-g^{*}$ and $\Delta^{\max}=\max_{j}\|b_{j}-\beta_{j}\|_{2}$. Define $\pi^{b,g}\left(X\right)=\arg\min_{j}b^{\top}X_{j}-g_{j}$, and let
@@
Let Assumption 3.1 and 3.2 hold. Let $T$ be a deterministic subset of $\left\{1,...,n\right\}$, and define $V_{T}=\sum_{n\in T}X^{n}\left(X^{n}\right)^{\top}$. Then, for any $z>0$,
@@
Let Assumptions 3.1 and 3.2 hold. Let $\left\{\pi^{n}\right\}^{\infty}_{n=1}$ be a sequence of random variables such that each $\pi^{n}$ takes values in $\left\{1,...,K\right\}$ and is measurable with respect to $\bar{\mathcal{F}}^{n}$, the sigma-algebra generated by $\pi^{1},...,\pi^{n-1}$, $W^{1}_{\pi^{1}},...,W^{n-1}_{\pi^{n-1}}$, and $X^{1},...,X^{n}$.
@@
Let Assumptions 3.1 and 3.2 hold, and let $\pi$ be the semi-myopic policy of Section 3.2. There exist constants $C_{3},C_{4},C_{5}>0$, which depend only on $d$, $K$ and $\kappa_{4},\kappa_{5},\kappa_{6}$, such that
@@
Let Assumptions 3.1 and 3.2 hold. Suppose that the semi-myopic policy of Section 3.2 is used to sample cost functions, while $g^{n}$ is updated as in Algorithm 1. Let $C_{1}$ be the constant obtained from Lemma 5.5. There exists a constant $C_{6}>0$ such that
@@
Let Assumptions 3.1 and 3.2 hold, and let $\pi$ be the semi-myopic policy of Section 3.2. There exists a constant $C_{7}>0$ such that
@@
Let Assumptions 3.1 and 3.2 hold. Suppose that the semi-myopic policy of Section 3.2 is used to sample cost functions, while $g^{n}$ is updated as in Algorithm 1. There exists some $\bar{\alpha}>0$ such that
@@
$\leq$ $\|\delta^{n}\|^{2}_{2}+\frac{2\alpha^{2}}{\left(n+1\right)^{2}}-\frac{2\alpha}{n+1}\left(\zeta^{n+1}-p^{\hat{\beta}^{n},g^{n}}\right)^{\top}\delta^{n}-\frac{2\alpha D_{1}}{n+1}\min\left\{\|\delta\|^{2}_{2},\left(\kappa^{\prime}_{1}\right)^{2}\right\}$
@@
$\mathbb{E}\left(\|\delta^{n+1}\|^{2}_{2}\mid\mathcal{F}^{n}\right)$ $\leq$ $\|\delta^{n}\|^{2}_{2}+\frac{2\alpha^{2}}{\left(n+1\right)^{2}}-\frac{2\alpha D_{1}}{n+1}\min\left\{\|\delta\|^{2}_{2},\left(\kappa^{\prime}_{1}\right)^{2}\right\}$
@@
$\mathbb{E}\left(\|\delta^{n+1}\|^{2}_{2}\mid\mathcal{F}^{n}\right)$ $\leq$ $\left(1-\frac{2\alpha D_{1}}{n+1}\min\left\{1,\frac{\left(\kappa^{\prime}_{1}\right)^{2}}{D^{2}_{0}}\right\}\right)\|\delta^{n}\|^{2}_{2}+\frac{2\alpha^{2}}{\left(n+1\right)^{2}}$
@@
$\mathbb{E}\left(\|\delta^{n+1}\|^{2}_{2}\mid\mathcal{F}^{n}\right)$ $\leq$ $\left(1-\frac{2\alpha D_{1}}{n+1}\min\left\{1,\frac{\left(\kappa^{\prime}_{1}\right)^{2}}{D^{2}_{0}}\right\}\right)\|\delta^{n}\|^{2}_{2}+\frac{2\alpha^{2}}{\left(n+1\right)^{2}}$ (43)
@@
$\leq$ $\left(1-\frac{2\alpha D_{1}}{n+1}\min\left\{1,\frac{\left(\kappa^{\prime}_{1}\right)^{2}}{D^{2}_{0}}\right\}\right)\|\delta^{n}\|^{2}_{2}+\frac{2\alpha^{2}}{\left(n+1\right)^{2}}$
@@
$\mathbb{E}\left(\|\delta^{n+1}\|^{2}_{2}\right)$ $\leq$ $\left(1-\frac{2\alpha D_{1}}{n+1}\min\left\{1,\frac{\left(\kappa^{\prime}_{1}\right)^{2}}{D^{2}_{0}}\right\}\right)\mathbb{E}\left(\|\delta^{n}\|^{2}_{2}\right)+\frac{2\alpha^{2}}{\left(n+1\right)^{2}}$
@@
$\leq$ $\left(1-\frac{2\alpha D_{1}}{n+1}\min\left\{1,\frac{\left(\kappa^{\prime}_{1}\right)^{2}}{D^{2}_{0}}\right\}\right)\mathbb{E}\left(\|\delta^{n}\|^{2}_{2}\right)+\frac{2\alpha^{2}}{\left(n+1\right)^{2}}$
@@
$\leq$ $\left(1-\frac{2\alpha D_{1}}{n+1}\min\left\{1,\frac{\left(\kappa^{\prime}_{1}\right)^{2}}{D^{2}_{0}}\right\}\right)\mathbb{E}\left(\|\delta^{n}\|^{2}_{2}\right)+\frac{2\alpha^{2}}{\left(n+1\right)^{2}}$
@@ 6 Numerical illustrations
We illustrate our framework on two numerical examples with linear costs. Section 6.1 considers synthetic test instances with $X$ generated uniformly on the unit sphere. Section 6.2 applies our framework to learn facility locations in a geographical partitioning problem (a common application area of semidiscrete optimal transport). Both types of experiments seek to explore the performance of the semi-myopic policy in settings that lie outside the assumptions used in our theoretical analysis.
@@ 6.1 Synthetic linear costs
We consider a set of $1000$ synthetic test problems. In all of these problems, the data $X$ are generated uniformly on the unit sphere in $d=10$ dimensions; since $X$ does not have a density in $\mathbb{R}^{d}$, Assumption 3.1 does not hold. Each problem uses $K=5$ different vectors $\beta_{k}$, also generated uniformly on the unit sphere, and a different $p$ generated uniformly on the $d$-dimensional simplex.
@@
For ease of implementation, we modified the semi-myopic policy to conduct forced exploration probabilistically: that is, in the $n$th iteration, we select $k$ uniformly at random with probability $\frac{\left(\frac{1}{a}\log n\right)^{9}}{n}$, where we used $a=4.5$. The rest of the time, we select a cost function according to the myopic policy $\hat{\pi}^{n}$. We set $\rho^{n}\equiv 0.001$ for simplicity, and in order to enable use of recursive least squares. At each $n$, we evaluate the policy using the correct selection indicator $1_{\left\{\hat{\pi}^{n}=\pi^{*}\left(X^{n}\right)\right\}}$. By averaging this quantity across $10$ runs, each containing $1000$ iterations and starting from $g^{0}\equiv 0$, $\hat{\beta}^{0}_{k}\equiv 0$, we obtain a trajectory of the probability $P\left(\hat{\pi}^{n}=\pi^{*}\left(X^{n}\right)\right)$ of correct selection. We then average this trajectory over $1000$ test problems.
@@
The optimal policy $\pi^{*}$ for each test problem is precomputed via brute force by running (12) with the true cost coefficients for a very large number of iterations. As a benchmark, we also implement a policy that knows the true costs (and, therefore, does not need to conduct any exploration or collect any samples), but does not know $g^{*}$, and runs (12) in an online manner. We also calculate the average probability of correct selection for this policy, which allows us to evaluate any loss incurred by having to learn the coefficients.
(a) Sensitivity analysis with respect to noise.
(b) Sensitivity analysis with respect to $\alpha$.
@@
Figure 1: Average performance of semi-myopic policy under Gaussian noise.
@@
Figure 1(a) shows average performance for the semi-myopic policy with stepsize parameter $\alpha=5$ and two levels of Gaussian noise: $Var\left(\varepsilon\right)=0.02^{2}$ and $Var\left(\varepsilon\right)=0.2^{2}$. Since the benchmark policy knows the costs, it is unaffected by the noise variance. We see that, when the variance is low, there is virtually no performance loss caused by unknown costs. When the variance is high, the average PCS of the semi-myopic policy lags behind that of the known-cost policy by $0.039$ (and has narrowed this gap relative to the earlier iterations). Performance also does not appear to be very sensitive to the frequency of forced exploration: we also tested a much higher value $a=6$, for which the probability of forced exploration is only $0.0036$ by the end of the time horizon, but obtained very similar trajectories. This also indicates that, in an online implementation, we could allow forced exploration to happen quite infrequently while maintaining good performance.
Figure 1(b) uses the higher variance level and varies the stepsize parameter $\alpha$ over several orders of magnitude. We observe some sensitivity with respect to this parameter, but, interestingly, this sensitivity is partially smoothed out by the presence of noise: under known costs, the difference between $\alpha=0.5$ and $\alpha=50$ is far more pronounced than under unknown costs. By the end of the time horizon, there is virtually no difference between the two best-performing choices, and the third-best is relatively close as well. Only the largest choice $\alpha=500$ lags behind the others, perhaps reflecting the fact that the multiplicative constant in Theorem 3.1 is increasing in $\alpha$.
Figure 2(a) repeats the analysis of Figure 1(b) for the same experimental setup, with the sole difference that the noise follows a Laplace (rather than Gaussian) distribution with mean zero and $Var\left(\varepsilon\right)=0.2^{2}$. Thus, this experiment considers a situation where the assumption of sub-Gaussianity is violated (instead, the noise is sub-exponential). However, despite the heavier tails, the results are virtually identical to the normal case, indicating that our approach is fairly robust to violations of this assumption. We also compared noise levels similarly to Figure 1(a), but the results were again nearly identical to the Gaussian case and so we omit them here.
(a) Laplace noise.
(b) Binary observations.
Figure 2: Average performance of semi-myopic policy in non-normal settings.
Finally, Figure 2(b) uses the same setup for $X$ and $\beta_{k}$, but now we assume that each observation $W^{n+1}_{k}$ is generated from a Bernoulli distribution with success probability $\chi\left(\beta^{\top}_{k}X^{n+1}\right)=\frac{1}{1+e^{-\beta^{\top}_{k}X^{n+1}}}$. In other words, the true cost structure is an instance of logistic regression with coefficients $\beta_{k}$. However, we do not use logistic regression to estimate the coefficients, due to the high computational cost of estimating a GLM. Instead, we apply recursive least squares directly to the binary observations, treating the error $W^{n+1}_{k}-\chi\left(\left(\hat{\beta}^{n}_{k}\right)^{\top}X^{n+1}\right)$ as the residual. Thus, the semi-myopic policy works with a partially misspecified statistical model, though Chen & Ryzhov (2020) showed that recursive least squares is still consistent in this situation (i.e., we will still learn the true $\beta_{k}$ values given enough time). Since estimation is more difficult with misspecification, we extend the time horizon to $10^{4}$ iterations for this experiment.
Because the benchmark policy knows the costs, misspecification is not a concern for it, and its performance only depends on $\alpha$. The semi-myopic policy has much more difficulty in the misspecified setting than in the previous two, but makes consistent improvement throughout the majority of the time horizon. Because cost estimation is now the most difficult aspect of the problem, the performance of the semi-myopic policy is now almost entirely unaffected by $\alpha$.
It is well-known in the stochastic approximation literature that the performance of SA can be sensitive to the stepsize rule. It may be possible to speed up convergence by using other choices of $\alpha_{n}$, many of which are surveyed in George & Powell (2006). Other, more sophisticated procedures can be found in Broadie et al. (2011) or Schaul et al. (2013), but every one of these methods involves at least one tunable parameter, so there is no way to avoid the tuning issue entirely. However, it is noteworthy that the presence of noise in the observed costs has the effect of reducing sensitivity to $\alpha$ in all three of the settings considered here.